OpenStudy (anonymous):

PLEASE HELP f(x)=(x-11)^2 where x

zepdrix (zepdrix):

Hey Miss Britt c: Step 1: Rename your \(\Large\rm f(x)\) to \(\Large\rm y\). Step 2: Your \(\Large\rm x\)'s and \(\Large\rm y\)'s switch places. Step 3: Solve for your \(\Large\rm y\).

zepdrix (zepdrix):

\[\Large\rm f(x)=(x-11)^2\] Here is step 1:\[\Large\rm y=(x-11)^2\] Here is step 2:\[\Large\rm x=(y-11)^2\]

zepdrix (zepdrix):

From here, we want to isolate our \(\Large\rm y\). Hmm it's being squared... What can we do to "undo" the squaring? What is the inverse operation of squaring?

OpenStudy (anonymous):

square root

zepdrix (zepdrix):

Mmmm good good good, yah let's use that.

zepdrix (zepdrix):

\[\Large\rm \pm\sqrt{x}=y-11\]We usually end up with a plus/minus when taking the root of a square. I dunno if it's going to matter here though, mm thinking.

OpenStudy (anonymous):

cant be positive because of restriction

jimthompson5910 (jim_thompson5910):

if x <= 11, then y >= 0 for the original function f(x) the domain and range swaps because x and y have swapped. So the domain of the inverse is x >= 0 and the range for the inverse is y <= 11

zepdrix (zepdrix):

ah good point, both of you c: hehe

zepdrix (zepdrix):

\[\Large\rm \sqrt{x}=y-11\]

zepdrix (zepdrix):

After you've finished solving for y, replace it with the notation, \(\Large\rm f^{-1}(x)\).

jimthompson5910 (jim_thompson5910):

because the range is y <= 11, this means negative y outputs are possible. So you go for the negative form of the square root (not the positive)

zepdrix (zepdrix):

Woops, I forgot to write the new restriction >.< Sorry bout that.

zepdrix (zepdrix):

Ah i forgot to put the negative on the root as well :( Oh boy, it's not my night...

OpenStudy (anonymous):

oh thats what i did wrong, i forgot the negative sign too

zepdrix (zepdrix):

Ah :)

OpenStudy (anonymous):

hmm, how do i find the domain

OpenStudy (anonymous):

[11, inf) ?

zepdrix (zepdrix):

If you look back at what Jim was saying, Our `old range` becomes our `new domain`. Our range was, \(\Large\rm (0,\infty)\). So that becomes our domain.

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

thanks!

zepdrix (zepdrix):

Here is a graph of the functions, just in case you were curious :) https://www.desmos.com/calculator/hamwlcgtll The orange is the f(x), half of a parabola (because of the x restriction). Blue is the inverse function.